OpenStudy (anonymous):

The elements in this circuit have the following values: R1=100Ω, R2=100Ω, R3=200Ω, V1=5V, I1=0.01A and I2=0.001A. find va,vb,power of r1,power of r2,power of r3,power of v1? plz help me

OpenStudy (nurali):

Question 1 Formula vA = (R2*V1 + R3*V1 + I1*R1*R3 - I2*R1*R3)/(R1 + R2 + R3) vB = (R3*(V1 + I1*R1 + I1*R2 - I2*R1 - I2*R2))/(R1 + R2 + R3) P_R1 = (vA-V1)^2/R1 P_R2 = (vA-vB)^2/R2 P_R3 = vB^2/R3 P_V1 = V1 * (V1-vA)/R1 P_I1 = V1 * I1 P_I2 = V1* (-I2)

OpenStudy (nurali):

If correct please click best response

OpenStudy (anonymous):

For checking your answers ,,just draw circuit in lab & do dc analysis !!! you will get correct answers !!